# Kandinsky Conformal Prediction:

### v1

@@ Appendix B Proofs from Section 4
Let $\alpha$ be a parameter in $(0,1)$, and let ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, for any $w\in{\cal W}$, the prediction set given by Algorithm 3 satisfies
@@
$\mathbb{E}_{D,E}[\operatorname{wCD}({\cal C},\alpha,w)]$ $=\mathbb{E}\left[w(X_{n+1},Y_{n+1})\left(\mathbbm{1}\left\{Y_{n+1}\in{\cal C}(X_{n+1})\right\}-(1-\alpha)\right)\right]$
@@
We notice that $R\left(\left[\Phi(X_{j_{1}},Y_{j_{1}})|\ldots|\Phi(X_{j_{d+1}},Y_{j_{d+1}})\right]\right)$ is a $d$-dimensional subspace of $\mathbb{R}^{d+1}$. Since for fixed $\{\Phi(X_{i},Y_{i})\}_{i\in[n+1]}$ the scores ${\widetilde{S}}_{j_{1}},\ldots,{\widetilde{S}}_{j_{d+1}}$ are independent and continuously distributed, we have that for all $11\leq j_{1}<\ldots<j_{d+1}\leq n+$ 1
@@ C.1 ACSIncome
We preprocess the dataset following Liu et al. [2023]. We additionally apply logarithmic transformation of labels with base 10 and scale the label to $[0,1]$ by min-max scaling.
@@
We take the deterministic score function of Conformalized Quantile Regression (CQR). Given base predictors $f_{\alpha/2}(x)$ and $f_{1-\alpha/2}(x)$ for the $\alpha/2$ and $21-\alpha/$ 2 quantile, respectively,
@@ C.2 CivilComments
Following Koh et al. [2021], we split the dataset into 269,038 training samples and 178,962 samples for calibration and test. We finetune a DistilBERT-base-uncased model with a classification head on the training set, following the configurations of Koh et al. [2021]. We randomly redistribute samples between the calibration and test sets to vary the calibration sample sizes. Since the groups are overlapping, we estimate the ratio between the group average sample size and the overall sample number of 178,962. Then we downsample the dataset accordingly to approximately reach a prescribed group average sample size in the calibration set. We repeat the redistribution procedure 100 times, but the training set is fixed such that the DistilBert model is trained only once. However, we train the weight functions of Kandinsky conformal prediction on the calibration set, since we are considering the setup where training samples are only accessible to algorithms by the base predictor. Therefore, the weight function class is retrained for each calibration set. Since group sample sizes can be different between runs, in Figures 3(e) and 3(f), annotated group sample sizes are estimated by the mean of actual group sample sizes over 100 runs.

### v2

@@ Appendix B Proofs from Section 4
Let $\alpha$ be a parameter in (0,1)(0,1)( 0 , 1 ), and let ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. Then, for any $w\in{\cal W}$, the prediction set given by Algorithm 3 satisfies
@@
$\mathbb{E}_{D,E}[\operatorname{wCD}({\cal C},\alpha,w)]$ $=\mathbb{E}\left[w(X_{n+1},Y_{n+1})\left(\mathbbm{1}\left\{Y_{n+1}\in{\cal C}(X_{n+1};\varepsilon_{n+1})\right\}-(1-\alpha)\right)\right]$
@@
We notice that $R\left(\left[\Phi(X_{j_{1}},Y_{j_{1}})|\ldots|\Phi(X_{j_{d+1}},Y_{j_{d+1}})\right]\right)$ is a $d$-dimensional subspace of $\mathbb{R}^{d+1}$. Since for fixed $\{\Phi(X_{i},Y_{i})\}_{i\in[n+1]}$ the scores ${\widetilde{S}}_{j_{1}},\ldots,{\widetilde{S}}_{j_{d+1}}$ are independent and continuously distributed, we have that for all $1\leq j_{1}<\ldots<j_{d+1}\leq n+1$
@@ C.1 ACSIncome
We preprocess the dataset following Liu et al. [2023]. We additionally apply logarithmic transformation of labels with base 10 and scale the label to [0,1][0,1][ 0 , 1 ] by min-max scaling.
@@
We take the deterministic score function of Conformalized Quantile Regression (CQR). Given base predictors $f_{\alpha/2}(x)$ and $f_{1-\alpha/2}(x)$ for the $\alpha/2$ and $1-\alpha/2$ quantile, respectively,
@@ C.2 CivilComments
Following Koh et al. [2021], we split the dataset into 269,038 training samples and 178,962 samples for calibration and test. We finetune a DistilBERT-base-uncased model with a classification head on the training set, following the configurations of Koh et al. [2021]. We randomly redistribute samples between the calibration and test sets to vary the calibration sample sizes. Since the groups are overlapping, we estimate the ratio between the group average sample size and the overall sample number of 178,962. Then we downsample the dataset accordingly to approximately reach a prescribed group average sample size in the calibration set. We repeat the redistribution procedure 100 times, but the training set is fixed such that the DistilBert model is trained only once. However, we train the weight functions of Kandinsky conformal prediction on the calibration set, since we are considering the setup where training samples are only accessible to algorithms by the base predictor. Therefore, the weight function class is retrained for each calibration set. Since group sample sizes can be different between runs, in Figures 4(a) and 4(b), annotated group sample sizes are estimated by the mean of actual group sample sizes over 100 runs.
@@
C.3 MCQA
Kumar et al. [2023] takes the subset of MMLU benchmark containing 16 subjects spanning various fields and levels of study: professional accounting, computer security, high school computer science, college computer science, machine learning, formal logic, high school biology, anatomy, clinical knowledge, college medicine, professional medicine, college chemistry, marketing, public relations, management, and business ethics. They also evaluate the accuracy of the LLaMA-13B model on these subjects across ten different prompts (Figure 2, Kumar et al. [2023]). We further group the subjects into the basic and advanced category according to the medium accuracy of the model across all prompts. Nine subjects are classified as basic with medium accuracy over 40%: computer security, high school computer science, high school biology, anatomy, clinical knowledge, marketing, public relations, management, and business ethics. The remaining sevens subjects are classified as advanced: college computer science, machine learning, formal logic, college medicine, professional medicine, college chemistry, and professional accounting.